# Quantum Bayes Classifiers and Their Application in Image Classification

### v1

@@ Abstract
Bayesian networks are powerful tools for probabilistic analysis and have been widely used in machine learning and data science. Unlike the parameters learning mode of neural networks, Bayes classifiers only use sample features to determine the classification results without a time-consuming training process. We study the construction of quantum Bayes classifiers (QBCs) and design a naïve QBC and three semi-naïve QBCs (SN-QBCs). These QBCs are applied to image classification. A local features sampling method is employed to extract a limited number of features from images to reduce the computational complexity. These features are then used to construct Bayesian networks and generate QBCs. We simulate these QBCs on the MindQuantum quantum platform and test them on the MNIST and Fashion-MNIST datasets. Results show that these QBCs based on a limited number of features exhibit good classification accuracies. The classification accuracies of QBCs on the MNIST dataset surpass that of the classical Bayesian network and quantum neural networks that utilize all feature points.
@@ I Introduction
As a tool for studying causal relationships between variables and inferring the impact of variable states on outcomes, Bayesian networks are widely used in various fields such as machine learning and data science, including Monte Carlo analysis MosegaardSambridge-426 , reliability and risk analysis ModarresGroth-449 , health monitoring Daniels-420 , healthcare GaryJarrett-422 , and biomedical systems WongLin-428 . The size of a Bayesian network depends on the number of nodes and the dependencies among them Kwisthout-427 , and learning and inference in large networks can be challenging. Large-scale Bayesian networks have been proven to be an NP-hard problem ChickeringHeckerman-421 . The emergence of quantum computing has provided a new solution to this problem.
@@
In recent years, many quantum algorithms have demonstrated certain advantages over classical computation, providing acceleration in comparison. For example, the Shor’s algorithm Shor-97 achieves exponential acceleration compared to classical computation in solving the problem of factorizing large numbers. The Grover’s algorithm Grover-225 achieves quadratic acceleration in unstructured search. In addition, quantum algorithms based on classical machine learning, such as quantum support vector machine RebentrostMohseni-82 and quantum $K$-nearest neighbor DangJiang-419 , also demonstrate quantum acceleration properties. Image classification is a fundamental problem in computer vision. In recent years, several quantum machine learning models have been used for image classification. These models include quantum convolutional neural networks CongChoi-181 ; KerenidisLandman-432 ; HurKim-435 ; GongPei-455 , quantum $K$-nearest neighbor algorithm DangJiang-419 , and quantum ensemble methods SchuldPetruccione-326 ; MacalusoClissa-437 ; AraujoDaSilva-438 ; MacalusoLodi-431 ; ZhangWang-409 , etc.
@@
Quantum Bayesian networks (QBN) were introduced in 1995 as a simulation of classical ones Youssef-442 . In 2013, Ozols et al. proposed a quantum version of the rejection sampling algorithm called quantum rejection sampling for Bayesian inference OzolsRoetteler-445 . In 2016, Moreira and Wichert proposed a quantum-like Bayesian network that uses amplitudes to represent marginal and conditional probabilities MoreiraWichert-448 . In 2019, Woerner and Egger developed a quantum algorithm WoernerJ-446 for risk analysis using the principles of amplitude amplification and estimation. Their algorithm can provide a quadratic speed-up compared to classical Monte Carlo methods. In 2021, Borujeni et al. proposed a quantum circuit representation of Bayesian networks BorujeniNannapaneni-433 . They designed quantum Bayesian networks for specific problems, such as stock prediction and liquidity risk assessment.
@@
Different from the parameters learning mode of neural networks, Bayesian classification makes a classification decisions based only on sample features, without the tedious training process, resulting in lower computational complexity, faster speed, and less resource consumption. Currently, there is a lack of research on quantum Bayes classifiers (QBCs) building on Bayesian networks for image classification tasks. This paper studies QBCs for image classification. We design a naïve RishOthers-411 ; Leung-429 QBC and three semi-naïve QBCs (SN-QBCs), i.e., the SN-QBC based on SPODE network Kononenko-450 ; Zhou-423 with the feature in the center of an image as the superfather, the SN-QBC based on TAN network Kononenko-450 ; Zhou-423 , and the SN-QBC based on symmetric relationship of features in images. These QBCs are implemented using the MindQuantum Platform Mindquantum-418 , and the classification effects of QBCs are verified on the MNIST LecunBottou-417 and the Fashion-MNIST XiaoRasul-416 datasets.
@@
This paper is organized as follows. Sect. 2 introduces the basic knowledge of Bayes classifiers. Sect. 3 discusses the constructions of QBCs. Sect. 4 presents the image classification algorithm based on QBCs. Sect. 5 demonstrates the simulation results of QBCs for image classification on the MNIST and Fashion-MNIST datasets. Sect. 6 further discusses and concludes the paper.
@@ II Bayes classifier
A Bayes classifier is a statistical classifier based on Bayes’ theorem. It considers selecting the optimal category label based on probabilities and misclassification losses, assuming that all relevant probabilities are known. Suppose the feature of a sample data is $X=\{x_{1},x_{2},\cdots,x_{n}\}$, and the set of class labels is $Y=\{y_{1},y_{2},\cdots,x_{N}\}$. Based on the posterior probability $P(y_{i}|X)$, the expected loss of classifying the sample with feature $X$ as $y_{i}$ is defined as Zhou-423
@@
For a specific problem, the misclassification loss $\lambda_{ij}$ can be written as
@@
The optimal Bayes classifier selects the class that maximizes the posterior probability $P(y|X)$ given the sample $X$.
Obtaining an accurate posterior probability $P(y|X)$ is important, but in reality, this value is often challenging to obtain. In the probability framework, the posterior probability $P(y|X)$ can be estimated based on a finite training sample. According to Bayes’ theorem Zhou-423 , the posterior probability $P(y|X)$ can be written as
$P(y|X)=\frac{P(X|y)P(y)}{P(X)},$ (5)
where $P(X)$ is the evidence factor used for normalization, $P(y)$ is the class-prior probability, and $P(X|y)$ is the class-conditional probability of the sample $X$ with respect to $y$. According to the law of large numbers Offmann-RgensenPisier-451 , when there are a sufficient number of independently and identically distributed samples, $P(y)$ can be estimated by the frequency of each class that appears in the sample. As for the class-conditional probability $P(X|y)$, it involves combinations of all features in $X$. Assuming each feature $x_{i}$ has $d$ possible values, $P(X|y)$ will have $N*d^{n}$ possible values.
@@ II.1 Naïve Bayes Classifier
Clearly, it is difficult to directly obtain the class-conditional probabilities from a limited training sample, as it will result in the problem of combinatorial explosion in calculation, which becomes more severe with the increase of features. The naïve Bayes classifier RishOthers-411 ; Leung-429 is based on the assumption of “independence”, which assumes that each attribute independently affects the classification result, as shown in Fig. (1a). In this case, Eq. (5) can be rewritten as
@@ II.2 Semi-naïve Bayes Classifier
The premise of the naïve Bayes classifier is that the features satisfy the assumption of independence, but this is often not the case in practical applications. Therefore, the learning method of the semi-naïve Bayes classifier Kononenko-450 ; Zhou-423 has emerged. It properly considers the dependencies between certain features, which helps to avoid the problem of combinatorial explosion caused by considering the joint probability distribution. Additionally, it takes into account stronger feature dependencies. One-Dependent Estimator (ODE) is the most common strategy for the semi-naïve Bayes classifier, which assumes that each feature only depends on at most one other feature besides the label $y$, that is,
$P(y|X)\propto P(y)\prod_{i}^{n}P(x_{i}|y,pa_{i}),$ (8)
where $pa_{i}$ is the dependent feature (or parent feature) of $x_{i}$. The key problem of the semi-naïve Bayes classifier is how to determine the parent feature of feature $x_{i}$.
A typical approach is to assume that all features depend on the same feature, known as the Super-Parent ODE (SPODE), whose dependency relationship is shown in Fig. (1b). Another approach is the Tree Augmented Naïve Bayes (TAN), which computes the conditional mutual information between any pair of features and constructs a maximum weighted tree Camerini-452 ; Zhou-423 based on the feature dependencies, as shown in Fig. (1c).
@@
Figure 1: Bayesian networks.
@@ III Quantum Bayes classifiers
Unlike classical bits that can only represent either 0 or 1, a quantum bit (qubit) can represent both 0 and 1 simultaneously, i.e., quantum superposition. A single qubit state can be represented as |$\ket{\varphi}=\alpha\ket{0}+\beta\ket{1}$, where |$\ket{0}$ and |$\ket{1}$ are the basis states of a single qubit, $\alpha$ and $\beta$ are the amplitudes that satisfy the normalization condition $|\alpha|^{2}+|\beta|^{2}=1$. When |$\ket{\varphi}$ is measured in the computational basis {|$\{\ket{0},\ket{1}\}$, the state will collapse into basis states |$\ket{0}$ or |$\ket{1}$ with the probability $|\alpha|^{2}$ or $|\beta|^{2}$, respectively.
In the quantum gate computing model NielsenChuang-404 , quantum gates are used to represent unitary operations acting on qubits. Quantum gates can be divided into single-qubit gates and multi-qubit gates, and any multi-qubit gate can be decomposed into a set of universal quantum gates NielsenChuang-404 .
In this paper, a quantum circuit for a QBC is constructed using single qubit gates $R_{y}$ and $X$, and multi-qubit controlled gates $C^{n}R_{y}$. The $X$-gate is a flip gate that flips |$\ket{0}$ to |$\ket{1}$ or |$\ket{1}$ to |$\ket{0}$. Its matrix form is
@@
$R_{y}$ is a single qubit rotation gate which has the form
@@ III.1 Naïve-QBC
A QBC uses Bayes’ rule to perform classification tasks within the framework of quantum computing. A single qubit can be used to represent a binary node in a Bayesian network, and then a superposition quantum state can be used to represent the probability of different label values under various combinations of features in the Bayesian network, i.e., $P(y|X)$.
In a naïve Bayesian network, all features are independent of each other, and each feature only depends on the label feature. Let $n$ feature nodes be $x_{1}$, $x_{2}$, $\cdots$, $x_{n}$, and each feature depends only on the label feature $y$. The quantum circuit of the naïve quantum Bayesian classifier can be created as follows. The classifier circuit is composed of $n+1$ qubits, corresponding to the label feature $y$ and $n$ feature nodes $x_{1}$ to $x_{n}$, and all nodes are initialized to |$\ket{0}$.
For the label node $y$, the class prior probability $P(y=0)$ can be obtained by statistically counting enough independent sample data. After that, one can realize the quantum circuit representation of the class prior probabilities by $R_{y}(\cdot)$ operation. Unlike the encoding function $\arctan$ used in the Ref. BorujeniNannapaneni-433 , we use a more natural way of encoding, that is, let
@@
This achieves quantum encoding of the class-prior probabilities. For convenience in the following content, let
@@
While for each feature node $x_{i}$, one can get the class-conditional probabilities by statistically calculating $P(x_{i}|y=0)$ and $P(x_{i}|y=1)$. Then, these probabilities can be encoded by $CR_{y}(\cdot)$, where the controlled rotation angles are set as $\theta_{i+1}=f(P(x_{i}|y=0))$ and $\theta_{i+n+1}=f(P(x_{i}|y=1))$, respectively.
Figure 2: The quantum circuit for the dependency relationship of $y\rightarrow x_{1}$.
Take the naïve Bayesian network in Fig. 1a as an example. These are one label node $y$ and one feature node $x_{1}$, and each node has only 2 values, either 0 or 1. By statistically counting the class-prior probabilities $P(y=0)$ and the class-conditional probabilities $P(x_{1}|y=0)$ and $P(x_{1}|y=1)$, one can construct a quantum circuit as depicted in Fig. 2. In this case, the output of the circuit is
@@
Note that the probabilities of different label values and feature values are encoded in the amplitude of the output state, which is the same as Eq. (7). That is, the quantum circuit implements the naïve QBC. By continuously adding new feature nodes to the quantum circuit and establishing controlled rotations between parent nodes and child nodes, one can construct QBCs based on different dependency relationships. Another example is a Bayesian network with multiple features. That is, the naïve QBC with $n=3$ features is shown in Fig. 3a.
(a) The naïve QBC with 3 features.
(b) The SN-QBC based on SPODE structure with 3 features and $x_{1}$ as the superparent node.
Figure 3: Quantum circuits of QBCs.
In the prediction stage, for a given feature value $X=X^{\ast}$, one only needs to obtain the probabilities of the basis state |$\ket{y=i,X=X^{\ast}}$ by measurement. That is, obtain the values of $P(y=0,X=X^{\ast})$ and $P(y=1,X=X^{\ast})$, and choose the $y$ value associated with the higher probability as the classification outcome of the QBC.

### v2

@@ Abstract
Bayesian networks are powerful tools for probabilistic analysis and have been widely used in machine learning and data science. Unlike the time-consuming parameter training process of neural networks, Bayes classifiers constructed on Bayesian networks can make decisions based solely on statistical data from samples. In this paper, we focus on constructing quantum Bayes classifiers (QBCs). We design both a naïve QBC and three semi-naïve QBCs (SN-QBCs). These QBCs are then applied to image classification tasks. To reduce computational complexity, we employ a local feature sampling method to extract a limited number of feature attributes from an image. These attributes serve as nodes of the Bayesian networks to generate the QBCs. We simulate these QBCs on the MindQuantum platform and evaluate their performance on the MNIST and Fashion-MNIST datasets. Our results demonstrate that these QBCs achieve good classification accuracies even with a limited number of attributes. The classification accuracies of QBCs on the MNIST dataset surpass those of classical Bayesian networks and quantum neural networks that utilize all available feature attributes.
@@ I Introduction
As a powerful tool for studying causal relationships between variables and inferring the impact of variable states on outcomes, Bayesian networks are widely used in machine learning and data science, including Monte Carlo analysis MosegaardSambridge-426 , reliability and risk analysis ModarresGroth-449 , health monitoring Daniels-420 , healthcare GaryJarrett-422 , and biomedical systems WongLin-428 , etc. The size of a Bayesian network depends on the number of nodes and their dependencies Kwisthout-427 . Learning and inference in complex networks can be challenging, especially when dealing with large-scale Bayesian networks, which have been proven to be an NP-hard problem ChickeringHeckerman-421 . The emergence of quantum computing offers a new solution to this challenge.
@@
In recent years, many quantum algorithms have demonstrated quantum supremacy for achieving certain accelerations over their classical counterparts. For example, the Shor’s algorithm Shor-97 achieves exponential acceleration in solving the large number factorization problem. The Grover’s algorithm Grover-225 achieves quadratic acceleration in searching the unstructured data. In addition, quantum algorithms based on classical machine learning, such as quantum support vector machine RebentrostMohseni-82 and quantum $K$-nearest neighbor DangJiang-419 , have also demonstrated quantum accelerations. Image classification is a fundamental problem in computer vision. With the advancement of quantum machine learning, several quantum classifiers have been developed for image classification, including quantum convolutional neural networks CongChoi-181 ; KerenidisLandman-432 ; HurKim-435 ; GongPei-455 , quantum $K$-nearest neighbor algorithm DangJiang-419 , and quantum ensemble methods SchuldPetruccione-326 ; MacalusoClissa-437 ; AraujoDaSilva-438 ; MacalusoLodi-431 ; ZhangWang-409 , etc. Recent progresses of quantum classifiers could be found in LiDeng-479 .
@@
Quantum Bayesian networks (QBN) were introduced in 1995 as a simulation of classical ones Youssef-442 . In 2013, Ozols et al. proposed a quantum version of the rejection sampling algorithm called quantum rejection sampling for Bayesian inference OzolsRoetteler-445 . In 2016, Moreira and Wichert proposed a quantum-like Bayesian network that uses amplitudes to represent marginal and conditional probabilities MoreiraWichert-448 . In 2019, Woerner and Egger developed a quantum algorithm WoernerJ-446 for risk analysis using the principles of amplitude amplification and estimation. Their algorithm can provide a quadratic speed-up compared to classical Monte Carlo methods. In 2023, Gabriel et al. proposed a quantum algorithm based on quantum walks for quantum Bayesian estimation of gravitational waves parameters from black holes EscrigCampos-480 . Recently, Borujeni et al. proposed a quantum circuit representation of Bayesian networks BorujeniNannapaneni-433 . They designed quantum Bayesian networks for specific problems, such as stock prediction and liquidity risk assessment. Walid et al. further proposed an optimized version of quantum circuit for improving the quantum representation of Bayesian networks FathallahAmor-502 .
@@
Currently, there is a lack of research on quantum Bayes classifiers (QBCs) building on Bayesian networks for solving image classification problems. Different from the parameters learning mode of neural networks, a Bayes classifier makes a classification decisions based only on sample features, without the tedious training process, resulting in lower computational complexity, faster speed, and less resource consumption. In this paper, we study the construction of quantum Bayes classifiers (QBCs). Specifically, we design a naïve RishOthers-411 ; Leung-429 QBC and three semi-naïve QBCs (SN-QBCs), i.e., the SN-QBC based on SPODE network Kononenko-450 ; Zhou-423 with the attribute in the center of an image as the superfather, the SN-QBC based on TAN network Kononenko-450 ; Zhou-423 , and the SN-QBC based on symmetric relationship of attributes in an image. These QBCs are then applied to image classification tasks. We simulate these QBCs on the MindQuantum platform Mindquantum-418 and evaluate their performance on the MNIST LecunBottou-417 and Fashion-MNIST XiaoRasul-416 datasets.
@@
This paper is organized as follows. Sect. 2 introduces the basic concepts of Bayes classifiers. Sect. 3 discusses the constructions of QBCs. Sect. 4 presents the image classification algorithm based on QBCs. Sect. 5 demonstrates the simulation results of QBCs for image classification on the MNIST and Fashion-MNIST datasets. Sect. 6 further discusses and concludes the paper.
@@ II Bayes classifier
A Bayes classifier is a statistical classifier based on Bayes’ theorem. It considers selecting the optimal category label based on probabilities and misclassification losses, assuming that all relevant probabilities are known. Suppose the feature of a sample data is $X=\{x_{1},x_{2},\cdots,x_{i},\cdots,x_{n}\}$ with $x_{i}$ is an attribute of $X$, and the set of class labels is $Y=\{y_{1},y_{2},\cdots,y_{i},\cdots,y_{N}\}$. Based on the posterior probability $P(y_{i}|X)$, the expected loss of classifying the sample with feature $X$ as $y_{i}$ is defined as Zhou-423
@@
For a specific problem to minimize the misclassification rate, $\lambda_{ij}$ can be written as
@@
That is, the optimal Bayes classifier selects the class that maximizes the posterior probability $P(y|X)$ given the sample with feature $X$.
Obtaining an accurate posterior probability $P(y|X)$ is critical for a Bayes classifier, but this is often challenging in reality. In the probability framework, the posterior probability $P(y|X)$ can be estimated based on a finite number of training sample. According to Bayes’ theorem Zhou-423 , the posterior probability $P(y|X)$ can be written as
$P(y|X)=\frac{P(X|y)P(y)}{P(X)}$ (5)
where $P(X)$ is the evidence factor used for normalization, $P(y)$ is the class-prior probability, and $P(X|y)$ is the class-conditional probability of the sample $X$ with respect to $y$. According to the law of large numbers Offmann-RgensenPisier-451 , when there are a sufficient number of independently and identically distributed samples, $P(y)$ can be estimated by the frequency of each class that appears in the training set. As for the class-conditional probability $P(X|y)$, it involves combinations of all attributes in $X$. Assuming each attribute $x_{i}$ has $d$ possible values, $P(X|y)$ will have $N*d^{n}$ possible values.
@@ II.1 Naïve Bayes Classifier
Clearly, it is difficult to obtain the class-conditional probabilities directly from a limited number of training sample, as it will result in the problem of combinatorial explosion in calculation, which becomes more severe with the increase of attributes. The naïve Bayes classifier RishOthers-411 ; Leung-429 is based on the assumption of “independence”, which assumes that each attribute independently affects the classification result, as shown in Fig. (1a). In this case, Eq. (5) can be rewritten as
@@ II.2 Semi-naïve Bayes Classifier
The premise of the naïve Bayes classifier is that all attributes satisfy the assumption of independence, but this is often not the case in practical applications. Therefore, the learning method of the semi-naïve Bayes classifier Kononenko-450 ; Zhou-423 has emerged, which considers stronger dependencies between attributes while avoids the problem of combinatorial explosion caused by considering the joint probability distribution of all attributes. One-Dependent Estimator (ODE) is the most common strategy for the semi-naïve Bayes classifier, which assumes that each attribute only depends on at most one other attribute besides the label $y$, i.e.,
$P(y|X)\propto P(y)\prod_{i}^{n}P(x_{i}|y,pa_{i}).$ (8)
where $pa_{i}$ is the dependent attribute (or parent attribute) of $x_{i}$. The key problem of the semi-naïve Bayes classifier is how to determine the parent attribute of $x_{i}$.
A typical approach is to assume that all attributes depend on a single attribute, referred to as the Super-Parent ODE (SPODE), whose dependency relationship is illustrated in Fig. (1b). An alternative method is the Tree Augmented Naïve Bayes (TAN), which calculates the conditional mutual information between any pair of attributes and constructs a maximum weighted tree Camerini-452 ; Zhou-423 based on the attribute dependencies, as depicted in Fig. (1c).
@@
Figure 1: Bayesian networks with different attributes dependencies.
@@ III Quantum Bayes classifiers
Unlike classical bits that can only represent either 0 or 1, a quantum bit (qubit) can represent both 0 and 1 simultaneously, i.e., quantum superposition. A single-qubit state can be represented as |$\ket{\varphi}=\alpha\ket{0}+\beta\ket{1}$, where |$\ket{0}$ and |$\ket{1}$ are the basis states of the single-qubit, $\alpha$ and $\beta$ are the amplitudes that satisfy the normalization condition $|\alpha|^{2}+|\beta|^{2}=1$. When |$\ket{\varphi}$ is measured in the computational basis {|$\{\ket{0},\ket{1}\}$, the state will collapse into basis states |$\ket{0}$ or |$\ket{1}$ with the probability $|\alpha|^{2}$ or $|\beta|^{2}$, respectively. In the quantum gate computing model NielsenChuang-404 , quantum gates are used to represent unitary operations acting on qubits. Quantum gates can be divided into single-qubit gates and multi-qubit gates, and any multi-qubit gate can be decomposed into a set of universal quantum gates NielsenChuang-404 .
In this paper, a quantum circuit for a QBC is constructed using single-qubit gates $R_{y}$ and $X$, and multi-qubit controlled gates $C^{n}R_{y}$. The $X$-gate is a flip gate that flips |$\ket{0}$ to |$\ket{1}$ or |$\ket{1}$ to |$\ket{0}$. Its matrix form is
@@
$R_{y}$ is a single-qubit rotation gate which has the form
@@ III.1 Naïve-QBC
A QBC uses Bayes’ rule to perform classification tasks within the framework of quantum computing. A single-qubit can be used to represent a node in a Bayesian network, and then a superposition quantum state can be used to represent the probability of different label values under various combinations of attributes in the Bayesian network, i.e., $P(y|X)$.
In a naïve Bayesian network, all attributes are independent of each other they depends only on the label. Let $n$ attribute nodes be $x_{1}$, $x_{2}$, $\cdots$, $x_{n}$, and each attribute depends only on the label $y$. The quantum circuit of the naïve-QBC is composed of $n+1$ qubits, corresponding to the label $y$ and $n$ attribute nodes $x_{1}$ to $x_{n}$. The naïve-QBC is constructed as follows:
  1. (a)
    All $n+1$ qubits are initialized to |$\ket{0}$.
  2. (b)
    For the label node $y$, the class-prior probability $P(y=0)$ can be obtained by statistically counting the training set. After that, one can encode the class-prior probabilities by $R_{y}(\cdot)$ operation. Unlike the encoding function $\arctan$ used in the Ref. BorujeniNannapaneni-433 , we use a more natural way of encoding. That is, let
@@
    This achieves quantum encoding of the class-prior probabilities. To simplify the representation in the following content, let
@@
  3. (c)
    For each attribute node $x_{i}$, one can get the class-conditional probabilities by statistically calculating $P(x_{i}=0|y=1)$ and $P(x_{i}=0|y=0)$. Then, these probabilities are encoded by $CR_{y}(\cdot)$, where the controlled rotation angles are set as $\theta_{i+1}=f(P(x_{i}=0|y=1))$ and $\theta_{i+n+1}=f(P(x_{i}=0|y=0))$, respectively.
Take the naïve Bayesian network in Fig. 1a as an example. Suppose these are one label node $y$ and one attribute node $x_{1}$ ($n=1$), and each node has only 2 values, either 0 or 1. By statistically counting the class-prior probabilities $P(y=0)$ and the class-conditional probabilities $P(x_{1}=0|y=1)$ and $P(x_{1}=0|y=0)$, one can construct a quantum circuit as depicted in Fig. 2. In this case, the output of the circuit is
@@
Note that the probabilities of different label values and attribute values are encoded in the amplitude of the output state, which is the same as Eq. (7). That is, the quantum circuit implements the naïve-QBC for features with only one attribute.
Figure 2: The quantum circuit for $y\rightarrow x_{1}$, i.e., the naïve-QBC with only one attribute ($n=1$).
By continuously adding new attribute nodes to the quantum circuit and establishing controlled rotations between parent nodes and child nodes, one can construct QBCs based on different dependency relationships. Another example is a Bayesian network with multiple attributes. That is, the naïve QBC with $n=3$ attributes is shown in Fig. 3.
Figure 3: The naïve-QBC with 3 attributes. Quantum circuits of QBCs.
In the prediction stage, for a given feature value $X^{\ast}$, one only needs to obtain the probabilities of the basis state |$\ket{y=i,X=X^{\ast}}$ by measuring the output state of the naïve QBC. For a binary classification problem, one obtains the values of $P(y=0,X=X^{\ast})$ and $P(y=1,X=X^{\ast})$, and chooses the $y$ value with a higher probability as the classification outcome of the QBC.